A transformer on a pole near a factory steps the voltage down from 3600 V (rms) to 120 V (rms). The transformer is to deliver 1400 kW to the factory at 85% efficiency. Find values for the following.

(a) the power delivered to the primary​

________ kW

(b) the current in the primary​

________ A

(c) the current in the secondary​

________ A
Answer #1

(a)
P₂ = η·P₁
P₁ = P₂/η
P₁ = 1400 kW/0.85
P₁ = 1647.1 kW

Power delivered to the primary, P₁ = 1647.1 kW

(b)
P₁ = U₁· I₁
P₁/U₁ = I₁
I₁ = 1647.1 kW/ 3600
I₁ = 457.53 Amp
Current in the primary​, I₁ = 457.53 Amp

(c)
P2 = U2· I2
P2 /U2 = I2
I2 = (1400 * 10^3) / 120
I2 = 11666.7 Amp
Current in the Secondary, I2 = 11666.7 Amp